A chemist measures the energy change Delta H durin
Answer #1

a)
since ΔH is positive, heat is absorbed
It is endothermic reaction

b)
yes, absorbed

c)
number of moles of H2O = mass/molar mass
= 94.8 / 18
= 5.27 mol

heat absorbed when 2 mol of H2O reacts = 572 KJ
for 5.27 mol, heat absorbed = (572*5.27)/2 = 1506 KJ
Answer: 1.51*10^3 KJ